# Uncertainty Quantification for Evaluating Machine Translation Bias

### v1

@@ F M S F S M O F O M S F S M O F O M F M S F S M O F O M S O
Table 3: anova results: single effects of bias cues (Feminine, Masculine, Subject and Object) on $-\mathcal{H}$ (s3e). Values correspond to effect coefficients (deviations from a reference group). Boldface indicates statistical significance (p < 0.05). The sign of the values indicates whether the presence of the variable increases (positive) or decreases (negative) the mean $\mathcal{H}$ of the group containing the given variable value. Reference group is N for all columns except: ‘no name’ for Names, ‘no default’ for Default M, ‘unambiguous’ for Ambiguity.
Semantic Entropy Scores Vary With Regard to Bias Cues.
The results of the second experiment presented in Table 3 show that most bias cues in the data have a significant effect on the variance of $-\mathcal{H}$ (s3e), indicating that the tested models exhibit various social and linguistic biases.333 $-\mathcal{H}$ (s3e) shows the strongest sensitivity to bias cues compared to $-\mathcal{H}$ (se) and $-\mathcal{H}$ (ge). We also experiment with unnormalised $\mathcal{H}$ scores, the results of which are less comparable across metrics, bias types, and models. The full results are presented in Appendix F. All trends observed for $-\mathcal{H}$ (s3e) are also present in $-\mathcal{H}$ (se) and $-\mathcal{H}$ (ge), as well as for unnormalised $\mathcal{H}$. The results corroborate previous findings. The high absolute coefficient values in the Names column indicate that person names have an effect on gender translation even when a disambiguating pronoun is present. This is in line with Saunders and Olsen (2023), who have shown that both pronouns and names induce gender bias and are often not sufficient for full disambiguation. Secondly, the fact that some Russian nouns have a default masculine grammatical gender regardless of the context (Komova, 2024) is reflected in significant decrease in gender diversity for sentences containing such nouns (negative coefficients in the Default M columns indicating lower $-\mathcal{H}$). Thirdly, we observe that masculine biases generally reduce $-\mathcal{H}$ (negative coefficients in the M columns), while feminine biases tend to increase it (positive coefficients in the F columns). This suggests a general default toward masculine translations in models, with outputs becoming more similar under masculine biases and more varied under feminine ones. This finding aligns with Kuzucu et al. (2025), who show that model uncertainty is typically higher for minority groups.
Translation Accuracy Affects the Bias–Entropy Relationship Differently Across Levels of Analysis.
Having validated the s3e metric in the first experiment, we investigate the results of $\Delta\mathcal{H}$ (s3e) as a bias metric. Table 4 illustrates that some models (Opus-MT-UK, deb-Opus-MT-UK, m2m100-UK, Opus-MT-RU, m2m100-RU) exhibit the desired negative $\Delta\mathcal{H}$. Surprisingly, this result suggests that when it comes to ambiguous instances, contrary to the unambiguous cases, the models which perform better on translation accuracy (namely models for Spanish and French) are not generally less gender biased (model rankings according to all metrics are available in Appendix J). This finding mirrors the results in the Ambiguity column in Table 3, where $-\mathcal{H}$ increases for Ukrainian and Russian (positive coefficients), but not for Spanish and French. Higher $-\mathcal{H}$ for Ambiguous items (or negative $\Delta\mathcal{H}$) is expected for an unbiased model.
Lang. Model Unamb Amb $\Delta\mathcal{H}$
ES Opus-MT 1.23 1.12 0.09
deb-Opus-MT 0.97 0.89 0.08
m2m100 1.79 1.45 0.19
FR Opus-MT 1.79 1.43 0.20
deb-Opus-MT 1.21 1.08 0.11
m2m100 3.22 2.78 0.14
UK Opus-MT 1.96 2.16 -0.10
deb-Opus-MT 1.98 2.15 -0.09
m2m100 2.05 2.28 -0.11
RU Opus-MT 1.56 1.68 -0.08
deb-Opus-MT 1.05 0.97 0.08
m2m100 1.83 2.29 -0.25
Table 4: Unambiguous and Ambiguous $\mathcal{H}$ (s3e)
In contrast, we observe the expected effect of debiasing on the Spanish and French models (lower $\Delta\mathcal{H}$ for deb-Opus-MT in Table 4), suggesting that models which perform better overall are more susceptible to debiasing. The impact of debiasing is also reflected in Table 3, as the effects are mostly smaller (lower absolute values of coefficients) in debiased models compared to their non-debiased counterparts across languages, confirming that debiasing is at least partially effective.
Figure 2: Violin plots of binned comet scores and $\mathcal{H}$ (s3e) on ambiguous and unambiguous inputs. Low, medium and high comet scores from left to right, evaluated with human translations, multi-reference for ambiguous items.
In Figure 2, results are grouped by comet score bins for a more fine-grained analysis at the instance level. For the models with negative $\Delta\mathcal{H}$ (s3e) scores (Ukrainian and Russian), $\Delta\mathcal{H}$ is typically most pronounced for the highest-accuracy translations (e.g. ambiguous scores for m2m100-RU in Bin 3 are substantially higher than B1). Although debiasing does not reduce the overall $\Delta\mathcal{H}$ score for Ukrainian (see Table 4), it results in the largest improvement in the highest-quality translations: in bin B3, ambiguous $\mathcal{H}$ scores for deb-Opus-MT-UK are notably higher than those of the original model. This improvement is further supported by a substantial 8.41% drop in masculine focus noun inflections for Ukrainian, compared to 0.88–2.49% for other languages. We hypothesise that this is due to the limited training data in Ukrainian, which may lead to a less stable model that performs worse overall but is more responsive to debiasing in high-quality outputs. The relationship between translation accuracy and bias under ambiguity appears to differ depending on the level of analysis: between models, higher accuracy does not imply lower bias on ambiguous instances, whereas within models, higher-accuracy instances tend to show lower bias.
Qualitative Analysis
A qualitative analysis of the example in Table 1, presented with corresponding $\mathcal{H}$ values in Table 5, corroborates the quantitative findings in Figure 2. When translating the sentence in row 1, across all target languages, Opus-MT models consistently produce only the masculine variants of the focus noun (‘El mecánico’, ‘Le mécanicien’, ‘Механик’ and ‘Механiк’). In the anti-stereotypical case (row 2), all languages except Russian include both masculine and feminine forms (‘La mecánica’, ‘La mécanicienne’ and ‘Механiка’), indicating that these models are sensitive to the masculine stereotype even when the referent in the context is clearly feminine. For Russian, the models fail to generate any feminine constructions, even when the context is unambiguously feminine. This difference is evident in the $\mathcal{H}$ (s3e) scores, which are higher in row 2 than row 1 for the first three languages. Moreover, in the ambiguous case (row 3), all Opus-MT models produce only masculine nouns, regardless of language. Consequently, $\mathcal{H}$ (s3e) scores are generally higher for the unambiguous cases (mean of rows 1 and 2) than for the ambiguous case (row 3), except for Russian, where $\mathcal{H}$ remains low across all conditions. These observations are consistent with expectations for biased models: they default to stereotypical gender realisations when the pronoun is ambiguous and sometimes even when the context clearly suggests an anti-stereotypical interpretation.
7 Conclusion
Limitations
Ethics Statement
Acknowledgments
Ieva Raminta Staliūnaitė is supported by Huawei. Andreas Vlachos is supported by the ERC grant AVeriTeC (GA 865958). Julius Cheng is supported by a research grant from Imminent. The translations for the datasets were funded by the same Imminent research grant and we would like to thank Andrii Nikolaiev, Mykhailo Khokhlovych, Sarah Mokadem and Daniel Carter for their work on the translations.
Appendix A Further UQ discussion
@@
Appendix B Annotation of Names
Table 6 presents the names used for expanding the WinoMT dataset to include common names for masculine and feminine genders in French, Spanish, Ukrainian and Russian.
@@ Appendix B Annotation of Names
F Carla Anne Anna Anna
M Gabriel Victor Ivan Ivan
@@ F Carla Anne Anna Anna
Appendix C Translation Guidelines and Translator Recruitment Details
@@
Appendix D Human Annotation Quality
@@
Appendix E Overall Model Performance
@@
Model Opus-MT deb-Opus-MT m2m100
Dataset es fr uk ru es fr uk ru es fr uk ru
newstest2012 84.52 82.21 84.47 82.22 71.25 71.97
newstest2013 85.28 83.45 85.24 83.44 72.84 72.57
newstest2014 85.01 87.44 85.00 87.44 74.63 72.29
wmttest2023 74.58 79.02 74.58 79.02 56.49 55.93
wmttest2024 66.99 71.64 67.00 71.63 49.20 47.63
mean 84.90 83.56 70.79 79.37 84.86 83.55 70.79 79.36 72.05 73.06 52.85 58.62
@@
Appendix F anova Results
@@ F M S F S M O F O M S F S M O F O M F M S F S M O F O M S O
Table 8: anova results: single effects of bias cues (Feminine, Masculine, Subject and Object) on $-\mathcal{H}$ (s3e), $-\mathcal{H}$ (se) and $-\mathcal{H}$ (ge). Values correspond to effect coefficients (deviations from a reference group). Boldface indicates statistical significance (p < 0.05). The sign of the values indicates whether the presence of the variable increases (positive) or decreases (negative) the mean $\mathcal{H}$ of the group containing the given variable value. Reference group is N for all columns except: ‘no name’ for Names, ‘no default’ for Default M, ‘unambiguous’ for Ambiguity.
@@
Appendix G Gender Accuracy
@@
Appendix H Quality Estimation with Human Translations
@@
Appendix I Log Probability and Surprisal Scores
@@
Language Model LogProb (Correct) LogProb (Incorrect) s3e $I$ (Correct) s3e $I$ (Incorrect) se $I$ (Correct) se $I$ (Incorrect) ge $I$ (Correct) ge $I$ (Incorrect) ES Opus-MT -149.7 -149.78 7.83 8.88 0.3 0.35 0.33 0.3 deb-Opus-MT -149.19 -149.01 8.08 9.16 0.29 0.31 0.35 0.33 m2m100 -226.61 -227.29 23.61 26.03 0.41 0.4 0.42 0.43 FR Opus-MT -197.1 -195.11 9.18 9.89 0.73 0.72 0.24 0.29 deb-Opus-MT -196.98 -195.09 9.18 9.85 0.48 0.52 0.26 0.33 m2m100 -283.91 -281.71 186.42 194.52 0.49 0.4 0.43 0.48 UK Opus-MT -161.98 -161.14 147.6 152.15 0.6 0.54 0.22 0.22 Opus-MT-debiased -161.46 -160.68 150.52 153.76 0.49 0.47 0.23 0.23 m2m100 -241.0 -241.49 204.72 211.9 0.28 0.24 0.23 0.25 RU Opus-MT -170.72 -170.9 32.14 33.15 0.38 0.43 0.08 0.19 deb-Opus-MT -170.58 -170.75 32.31 33.27 0.32 0.37 0.06 0.17 m2m100 -220.25 -220.78 218.11 219.06 0.45 0.4 0.16 0.3
@@
Appendix J Rankings by Different Metrics
Table 13 presents the rankings of models according to various metrics employed in this study.
comet Unambiguous comet All Gender Acc Delta S Delta H
deb-Opus-MT-RU Opus-MT-ES m2m100-ES m2m100-ES m2m100-RU
Opus-MT-RU deb-Opus-MT-ES deb-Opus-MT-ES deb-Opus-MT-ES m2m100-UK
m2m100-RU Opus-MT-FR Opus-MT-ES Opus-MT-ES Opus-MT-UK
m2m100-UK deb-Opus-MT-FR deb-Opus-MT-FR deb-Opus-MT-FR deb-Opus-MT-UK
Opus-MT-ES Opus-MT-RU Opus-MT-FR m2m100-FR Opus-MT-RU
deb-Opus-MT-ES deb-Opus-MT-RU m2m100-FR Opus-MT-FR deb-Opus-MT-RU
Opus-MT-UK m2m100-FR Opus-MT-RU deb-Opus-MT-RU deb-Opus-MT-ES
deb-Opus-MT-UK m2m100-ES m2m100-RU m2m100-RU Opus-MT-ES
Opus-MT-ES Opus-MT-UK deb-Opus-MT-RU deb-Opus-MT-UK deb-Opus-MT-FR
Opus-MT-FR deb-Opus-MT-UK m2m100-UK Opus-MT-UK m2m100-FR
deb-Opus-MT-FR m2m100-RU deb-Opus-MT-UK m2m100-UK m2m100-ES
m2m100-FR m2m100-UK Opus-MT-UK Opus-MT-RU Opus-MT-FR
Table 13: Model rankings across five evaluation metrics.
Appendix K Rank Correlation
@@
Spearman $\Delta I$ (s3e) -0.78 0.00
$\Delta I$ (se) -0.37 0.24
$\Delta I$ (ge) 0.27 0.00
ΔLog $\Delta\text{Log prob}$ 0.11 0.73
Kendall $\Delta I$ (s3e) -0.58 0.01
$\Delta I$ (se) -0.27 0.25
$\Delta I$ (ge) 0.23 0.00
ΔLog $\Delta\text{Log prob}$ 0.09 0.74
Table 14: Spearman and Kendall correlations between $\Delta I$ under different uncertainty metrics and Log Probabilities on the one hand, and gender accuracy on the other. Statistically significant correlations (p $<$ 0.05) are in bold.
Appendix L Entropy Scores
@@
Table 15: $\mathcal{H}$ scores across models and languages, with relative differences ($\Delta\mathcal{H}$) between unambiguous and ambiguous conditions.

### v2

@@ F M S F S M O F O M S F S M O F O M F M S F S M O F O M S O
Table 7: anova results: single effects of bias cues (Feminine, Masculine, Subject and Object) on norm-$\mathcal{H}$ (s3e). Values correspond to effect coefficients (deviations from a reference group). Boldface indicates statistical significance (p < 0.05). The sign of the values indicates whether the presence of the variable increases (positive) or decreases (negative) the mean $\mathcal{H}$ of the group containing the given variable value. Reference group is N for all columns except: ‘no name’ for Names, ‘no default’ for Default M, ‘unambiguous’ for Ambiguity.
Table 7 further demonstrates the impact of debiasing, with effect sizes typically reduced (smaller absolute coefficients) in debiased models versus non-debiased ones across languages, supporting the partial effectiveness of debiasing.
Appendix B Further UQ discussion
@@
Appendix C Annotation of Names
Table 8 presents the names used for expanding the WinoMT dataset to include commonly used person names for masculine and feminine genders in French, Spanish, Ukrainian and Russian.
@@ Appendix C Annotation of Names
F Carla, Lucía, María, Sofía, Ana Anne, Camille, Élise, Sophie, Claire Anna, Olena, Sofiia, Natalka, Oksana Anna, Maria, Elena, Svetlana, Natalia
M Gabriel, Carlos, Luis, Javier, Miguel Victor, Julien, Antoine, Lucas, Mathieu Ivan, Andrii, Taras, Oleh, Mykola Ivan, Dmitry, Alexei, Sergey, Nikolai
@@
Appendix D Translation Guidelines and Translator Recruitment Details
@@
Appendix E Human Annotation Quality
@@
Appendix F Overall Model Performance
@@
Dataset Model es fr uk ru de it el
newstest2012 Opus-MT 84.52 82.21
deb-Opus-MT 84.47 82.22
m2m100 84.23 82.69
Tower+9B 86.61 84.86
newstest2013 Opus-MT 85.28 83.45
deb-Opus-MT 85.24 83.44
m2m100 84.83 83.66
Tower+9B 87.14 86.27
newstest2014 Opus-MT 85.01 87.44
deb-Opus-MT 85.00 87.44
m2m100 85.28 71.95
Tower+9B 88.71 93.04
wmttest2023 Opus-MT 74.58 79.02
deb-Opus-MT 74.58 79.02
m2m100 73.58 80.23
Tower+9B 86.24 83.75
wmttest2024 Opus-MT 66.99 71.64
deb-Opus-MT 67.00 71.63
m2m100 79.45 89.30
Tower+9B 87.07 85.69
newstest2009 Opus-MT 81.30 84.65
deb-Opus-MT 81.24 84.58
m2m100 82.29 85.04
Tower+9B 85.33 87.78
newstest2024 Opus-MT 78.41
deb-Opus-MT 77.90
m2m100 78.58
Tower+9B 66.23
mean Opus-MT 84.90 83.56 70.79 79.37 81.30 84.65 78.41
deb-Opus-MT 84.86 83.55 70.79 79.36 81.24 84.58 77.90
m2m100 84.53 83.88 76.52 80.49 82.29 85.04 78.58
Tower+9B 86.88 86.61 86.66 87.49 85.33 87.78 66.23
@@
Appendix G anova Results
@@ F M S F S M O F O M S F S M O F O M F M S F S M O F O M S O
Table 10: anova results: single effects of bias cues (Feminine, Masculine, Subject and Object) on norm-$\mathcal{H}$ (s3e), norm-$\mathcal{H}$ (se) and norm-$\mathcal{H}$ (ge). Values correspond to effect coefficients (deviations from a reference group). Boldface indicates statistical significance (p < 0.05). The sign of the values indicates whether the presence of the variable increases (positive) or decreases (negative) the mean $\mathcal{H}$ of the group containing the given variable value. Reference group is N for all columns except: ‘no name’ for Names, ‘no default’ for Default M, ‘unambiguous’ for Ambiguity.
@@
Appendix H Gender Accuracy
@@
Appendix I Quality Estimation with Human Translations
@@
Appendix J Log Probability and Surprisal Scores
@@
Language Model LogProb (Correct) LogProb (Incorrect) s3e $I$ (Correct) s3e $I$ (Incorrect) se $I$ (Correct) se $I$ (Incorrect) ge $I$ (Correct) ge $I$ (Incorrect) ES Opus-MT -149.7 -149.78 7.83 8.88 0.3 0.35 0.33 0.3 deb-Opus-MT -149.19 -149.01 8.08 9.16 0.29 0.31 0.35 0.33 m2m100 -226.61 -227.29 23.61 26.03 0.41 0.4 0.42 0.43 Tower+9B -586.29 -586.72 15.25 19.80 0.97 1.19 0.51 0.51 FR Opus-MT -197.1 -195.11 9.18 9.89 0.73 0.72 0.24 0.29 deb-Opus-MT -196.98 -195.09 9.18 9.85 0.48 0.52 0.26 0.33 m2m100 -283.91 -281.71 186.42 194.52 0.49 0.4 0.43 0.48 Tower+9B -661.37 -677.52 15.62 19.79 1.24 1.33 0.78 0.78 UK Opus-MT -161.98 -161.14 147.6 152.15 0.6 0.54 0.22 0.22 Opus-MT-debiased -161.46 -160.68 150.52 153.76 0.49 0.47 0.23 0.23 m2m100 -241.0 -241.49 204.72 211.9 0.28 0.24 0.23 0.25 Tower+9B -455.91 -462.70 16.84 18.81 1.5 1.59 1.62 1.47 RU Opus-MT -170.72 -170.9 32.14 33.15 0.38 0.43 0.08 0.19 deb-Opus-MT -170.58 -170.75 32.31 33.27 0.32 0.37 0.06 0.17 m2m100 -220.25 -220.78 218.11 219.06 0.45 0.4 0.16 0.3 Tower+9B -499.57 -508.94 27.34 29.75 1.95 2.11 1.0 0.78
@@
Appendix K Results on the mGeNTe dataset
Table 15 presents the results of the experiments on the mGeNTe dataset.
Lang. Model Gender Acc $\Delta\text{Log prob}$ $\Delta I$ (s3e) comet ES Opus-MT 60.40 -0.06 -1.92 84.90 deb-Opus-MT 60.67 -0.06 -1.91 84.86 m2m100 60.53 -0.05 1.73 84.53 Tower+9B 67.47 -0.04 4.53 86.88 DE Opus-MT 69.20 -0.06 -1.45 81.30 deb-Opus-MT 68.93 -0.06 -1.43 81.24 m2m100 68.53 -0.06 1.96 82.29 Tower+9B 79.87 -0.04 2.21 85.33 IT Opus-MT 64.40 -0.04 -1.45 84.65 deb-Opus-MT 64.40 -0.04 -1.46 84.58 m2m100 66.00 -0.04 1.96 85.04 Tower+9B 74.00 -0.03 2.21 87.78 EL Opus-MT 65.73 -0.05 -1.43 78.41 deb-Opus-MT 65.87 -0.05 -1.41 77.90 m2m100 64.27 -0.08 1.96 78.58 Tower+9B 64.11 -0.02 2.18 66.23
Table 15: Gender Accuracy, $\Delta$ Log Probability, and $\Delta I$ (s3e) on Unambiguous instances of mGeNTe, comet scores on WMT test sets.
There is no correlation between $\Delta I$ scores and gender accuracy on the mGeNTe dataset (see Table 16).
Correlation Metric Statistic p-value
Spearman $\Delta I$ (s3e) -0.40 0.12
$\Delta I$ (se) -0.35 0.19
$\Delta I$ (ge) 0.44 0.09
$\Delta\text{Log prob}$ -0.11 0.70
Kendall $\Delta I$ (s3e) -0.28 0.14
$\Delta I$ (se) -0.18 0.34
$\Delta I$ (ge) 0.30 0.11
$\Delta\text{Log prob}$ -0.10 0.59
Table 16: Spearman and Kendall correlations between $\Delta I$ under different uncertainty metrics and Log Probabilities on the one hand, and gender accuracy on the other on the mGeNTe dataset. No statistically significant correlations present.
Table 17 presents the entropy (s3e) scores on the mGeNTe dataset. The best performing model (Tower+9B) shows the lowest $\Delta\mathcal{H}$ across languages, indicating the least bias.
Lang. Model Unamb Amb $\Delta\mathcal{H}$
ES Opus-MT 2.53 2.57 -0.02
deb-Opus-MT 2.45 2.47 -0.01
m2m100 2.41 2.28 0.06
Tower+9B 3.77 4.08 -0.08
DE Opus-MT 2.90 2.87 0.01
deb-Opus-MT 2.92 2.90 0.01
m2m100 2.72 2.90 -0.06
Tower+9B 3.36 4.35 -0.25
IT Opus-MT 2.89 2.78 0.04
deb-Opus-MT 2.17 1.79 0.19
m2m100 2.91 2.63 0.10
Tower+9B 3.77 4.08 -0.08
EL Opus-MT 2.80 2.67 0.05
deb-Opus-MT 2.75 2.60 0.05
m2m100 3.19 2.88 0.10
Tower+9B 6.74 6.99 -0.04
Table 17: Unambiguous and Ambiguous $\mathcal{H}$ (s3e) on mGeNTe. Statistically significant correlations between the $\Delta\mathcal{H}$ and comet scores at the instance level are presented in bold.
Appendix L Rankings by Different Metrics
Table 18 presents the rankings of models according to the metrics employed in this study.
Gender Acc Delta S
tower-ES tower-ES
m2m100-ES tower-FR
tower-FR deb-Opus-MT-ES
deb-Opus-MT-ES m2m100-ES
Opus-MT-ES tower-UK
deb-Opus-MT-FR Opus-MT-ES
Opus-MT-FR tower-RU
m2m100-FR deb-Opus-MT-FR
tower-UK m2m100-FR
tower-RU Opus-MT-FR
Opus-MT-UK Opus-MT-RU
m2m100-UK deb-Opus-MT-UK
deb-Opus-MT-UK deb-Opus-MT-RU
m2m100-RU m2m100-UK
deb-Opus-MT-RU m2m100-RU
Opus-MT-RU Opus-MT-UK
Table 18: Model rankings for gender accuracy (higher is better) and s3e delta (lower is better) on the WinoMT dataset.
Appendix M Rank Correlation
@@
Spearman $\Delta I$ (s3e) -0.84 0.00
$\Delta I$ (se) -0.42 0.10
$\Delta I$ (ge) 0.29 0.27
$\Delta\text{Log prob}$ 0.00 1.00
Kendall $\Delta I$ (s3e) -0.63 0.00
$\Delta I$ (se) -0.30 0.12
$\Delta I$ (ge) 0.21 0.27
$\Delta\text{Log prob}$ -0.03 0.89
Table 19: Spearman and Kendall correlations between $\Delta I$ under different uncertainty metrics and Log Probabilities on the one hand, and gender accuracy on the other, for the WinoMT dataset. Statistically significant correlations (p << 0.05) are in bold.
Appendix N Entropy Scores
@@
Table 20: $\mathcal{H}$ scores across models and languages, with relative differences ($\Delta\mathcal{H}$) between unambiguous and ambiguous conditions on the WinoMT dataset.